Project

General

Profile

2018-10-02-MEMOMIPO » History » Version 2

Long, Kenneth, 02 October 2018 17:34

1 1 Long, Kenneth
h2. Joint MICE Experiment Management Office and MICE International Project Office
2
3
h3. 18th September 2018; 15:30 BST
4
5
"Phone details":http://mice.iit.edu/phonebridge.html
6
7
h4. Agenda
8
9
# *Introduction: KL*
10
# *[[2018-09-18-MEMOMIPO|Notes of previous meeting]] and status of actions: All*
11
** *KL:* will migrate EB and CB mailing lists to jiscmail
12
** *CR:* will migrate analysis mailing list to jiscmail
13 2 Long, Kenneth
** *PF:* will migrate software mailing list to jiscmail
14 1 Long, Kenneth
# *Project Manager's report: CW*
15 2 Long, Kenneth
# *Software and computing status: PF*
16 1 Long, Kenneth
# *Paper status and plans: CR*
17
# *CM52: status of planning etc.: KL/All*
18
# *AOB*
19
20
h4. Notes
21
22 2 Long, Kenneth
*Present:* P. Soler, C. Whyte, C. Rogers, K. Long, A. Bross, 
23
24
# *Introduction: KL*
25
# *[[2018-09-18-MEMOMIPO|Notes of previous meeting]] and status of actions: All*
26
** *KL:* will migrate EB and CB mailing lists to jiscmail
27
*** Stands
28
** *CR:* will migrate analysis mailing list to jiscmail
29
*** Stands
30
** *PF:* will migrate software mailing list to jiscmail
31
*** Stands
32
# *Project Manager's report: CW*
33
** All detectors are now shipped.  Mississippi colleagues have removed the aerogel.  
34
** MICE Hall decommissioning continues: H2-gas panel and hydride bed has been removed.  South Mezzanine mods are underway.
35
** PPD ownership now for experiment.  Hole in synchrotron wall will be blocked on the inside so that there are no more any RP issues.
36
** Quads are being removed, most likely they will be scrapped unless a home can be found.  These quads need de-ionized water and have some earth-leakage issues.  *PS* will contact CERN East Area Test Beams to see if there is interest.
37
** Focus coils will be stored, being organised by J. Boehm.
38
** PS brought up issie of recovery of TDC purchased for MICE.  *CW* will organise recovery of boards for CM52.
39
** Organise a walk-around in the MICE Hall to see whether there is any equipment still worth recovering (action *CW*).
40
** Tracker Lab is to be repurposed; need to plan recovery of the equipment.  (action *KL/ABr*)
41
# *Software and computing status: PF*
42
** No Paolo.
43
# *Paper status and plans: CR*
44
** _Tracker analysis/reconstruction:_ Considering a noise-reduction algorithm based on FADC count or NPE.  This will then be used to inform improvements in pattern recognition.
45
** _Emittance-measurement paper:_ Revisit of consideration of apertures did not yield an apperture.  Now moving 'experiment' to correct analysis.  Indications are that this will help.
46
** _Emittance-evolution paper:_ Referees meeting tomorrow.  Awaiting result of this meeting.
47
** _System-performance paper:_ No update since meeting on paper last week.  CM52 we are expecting a draft of the paper.
48
** _6D/wedge paper:_ C. Brown, following LH2 analysis, will pick up this analysis.
49
** _Scattering:_ J.Nugent plans to update this week.  PS reported that following resolution of computing issues, JN is re-evaluating the alignment.  
50
** _LH2 scattering:_ G. Chapzitheodoridis is now progressing the LH2 scattering analysis.  LiH results have been reproduced and he intends to present his results at the CM.
51
** _Field-on scattering:_ A. Young continues to work on this.
52
** Students have been at RAL Summer School, so limitted progress in recent past.
53
** LH2 volume uniformity of density has been discussed with J. Boehm and C. MacWaters.  Mis-calibration may have been introduced by current limiters used to limit degree of heating produced by level sensors in LH2.
54
# *CM52: status of planning etc.: KL/All*
55
** *KL* will send email reminding people that the accommodation block booking is almost saturated and will be dropped tomorrow.
56
# *AOB*
57
** Pressure on office space is mounting.  *KL* will email collaboraiton to say need to clear so that we can reduce space use by factor of 2 on timescale of Christmas.
58
** D. Rajaram back for three months.  Working on track reconstruction with a view to making sure we dont have to look at it again when he is done.  We agree to invite him to MEMO/EB.  
59
60 1 Long, Kenneth
h4. Summary of actions:
61 2 Long, Kenneth
62
** *KL:* will migrate EB and CB mailing lists to jiscmail
63
** *CR:* will migrate analysis mailing list to jiscmail
64
** *PF:* will migrate software mailing list to jiscmail
65
** *CW:* organise recovery of TDC boards at CM52.
66
** *CW:* set up 'walk around' of Hall at CM52 to identify remaining valuable equipment.
67
** *KL/ABr:* Organise recovery of tracker equipment from tracker lab at CM52.
68
** *KL:* Remind collaboration of imperative to book accommodation for CM52 asap.
69
** *KL:* Inform collaboraiton of impending reduction in office space.
70
** *KL:* Invite DR to MEMO/EB.
71 1 Long, Kenneth
72
h3. Back to MEMO:
73
74
[[memo:|Back]]
75
76
h3. Back to MIPO:
77
78
[[mipo:|Back]]
79
80
h2. Documents